Solve for x
x-9=-x+a+4b

Answer by tara0066(31)

Solve for x
x-9=-x+a+4b First off, I would add x to both sides. This will cancel your (-x).
x(+x)-9=-x(+x)+a+4b Simplify
2x-9=a+4b Next add 9 to both sides to cancel out your (-9)
2x-9(+9)=a+4b(+9) Simplify
2x=a+4b+9 Then, divide both sides by 2. This will cancel the two before your x.
2x(/2)=(a+4b+9)(/2)
x=(a+4b+9)(/2)
Could also be written:
x=1/2a+2b+9/2 or x=1/2a+2b+4 1/2